# Entropy-Lens: The Information Signature of Transformer Computations

### v1

@@ Abstract
Transformer models have revolutionized fields from natural language processing to computer vision, yet their internal computational dynamics remain poorly understood—raising concerns about predictability and robustness. In this work, we introduce Entropy-Lens, a scalable, model-agnostic framework that leverages information theory to interpret frozen, off-the-shelf large-scale transformers. By quantifying the evolution of Shannon entropy within intermediate residual streams, our approach extracts computational signatures that distinguish model families, categorize task-specific prompts, and correlate with output accuracy. We further demonstrate the generality of our method by extending the analysis to vision transformers. Our results suggest that entropy-based metrics can serve as a principled tool for unveiling the inner workings of modern transformer architectures.
footnotetext: Equal contribution.
@@ 1 Introduction
Transformer-based architectures (Vaswani et al., 2023) are widely employed as state of the art models in several fields, from machine translation and search engines to DNA analysis and protein research (Devlin, 2018; Khattab and Zaharia, 2020; Ji et al., 2021; Chandra et al., 2023). Their declination in language modeling of large corpora is referred to as large language models (LLMs), and in computer vision as vision transformers (ViTs). Despite their success and ubiquity, transformers’ inner workings remain largely unknown, resulting in unpredictable behaviour (Wei et al., 2022) and reliability concerns (Schroeder and Wood-Doughty, 2025; Huang et al., 2025).
Therefore, considerate research efforts are devoted to transformer-based architectures interpretability, mostly focusing on LLMs (Nanda and Bloom, 2022; Bereska and Gavves, 2024; Elhage et al., 2021, 2022) and ViTs (Chefer et al., 2021). While exciting results have been achieved in this area, they remain limited to toy models and simplified setups, both very different from real use-case conditions. Moreover, these methods often require training a set of probes (nostalgebraist, 2020; Belrose et al., 2023) or full models on ad-hoc tasks (Nanda et al., 2023), making them architecture or even model specific and computationally expensive. These limitations restrict the scope and usability of current methodologies, rendering them unsuitable for off-the-shelf or large-scale transformer-based architectures.
Hence, we develop Entropy-Lens, a scalable framework to address these limitations. Our methodology is architecture agnostic and applicable to frozen off-the-shelf large-scale transformers. In our experiments, we consider both LLMs, including Llama (Touvron et al., 2023), Gemma (Team et al., 2024), GPT (Radford, 2018) up to 9B parameters, and ViTs (Wu et al., 2020) and data-efficient image transformers (DeiTs) (Touvron et al., 2021; Wightman, 2019). In particular, we analyze the evolution of the generated tokens’ Shannon entropy after each intermediate block in the residual stream, as described in Section 4. In Section 5, we show that these quantities constitute an information-theoretic signature of which and how the computation is performed. Finally, our work paves the way for several potential research directions involving information theory for LLM interpretability, as outlined in Section 6.
Our contributions are as follows:
  1. 1.
    We develop a scalable model agnostic methodology grounded in information theory for frozen off-the-shelf large-scale transformer architectures interpretability (Section 4).
  2. 2.
    We analyze the Shannon entropy of LLMs’ generated tokens’ intermediate predictions, showing that these quantities (1) identify the model family that generated them (2) identify the ‘task type’ of the prompt (3) correlate with the correctness of LLM generated answers to multiple choice questions (Section 5).
  3. 3.
    We demonstrate that the same framework is adaptable to domains outside of language modeling. In particular, we apply it to computer vision to show that the Shannon entropy of ViTs’ intermediate predictions identify the model family that generated them (Section 5), similarly to the LLMs case.
@@ 2 Related Work
Mechanistic interpretability (Bereska and Gavves, 2024) aims to provide a precise description and prediction of transformer-based computations. Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream. For example, logit-lens (nostalgebraist, 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space. tuned-lens (Belrose et al., 2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function. Building on the Transformer-Lens library (Nanda and Bloom, 2022), we propose Entropy-Lens, which employ logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
Transformers’ Circuits
Another approach to mechanistic interpretability aims to identify and understand the specific sub-computations, or circuits, in a neural network (Saphra and Wiegreffe, 2024). Olah et al. (2020) pioneered this approach, introducing the concept of circuits and demonstrating their existence in small models through manual analysis. In transformers, circuits are hypothesized to act as agents that read from and write to the residual stream, which acts as a form of memory. This has been demonstrated in a simplified transformer model composed only of attention blocks (without MLPs) (Elhage et al., 2021). However, there is evidence that full transformers exhibit similar behavior. This evidence—linked to the concept of superposition, the idea that models can represent more features than the available dimensions by compressing multiple features into one (Elhage et al., 2022)—is supported by studies on sparse autoencoders, which demonstrate the ability to decompose representations into simpler components (Bricken et al., 2023). Conmy et al. (2023a) developed a toolkit to facilitate mechanistic interpretability, offering techniques like activation patching and weight factorization. Building upon this, Conmy et al. (2023b) explored automated circuit discovery methods, addressing the challenge of scaling analysis to larger models. These works collectively emphasize the importance of understanding the concrete computational steps within LLMs, moving beyond superficial observations to reveal the underlying mechanisms.
@@
Information theory has been studied both in connection to the training phase of LLMs and their interpretability. For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al., 2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al., 2024). Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton et al., 2024). Our work, instead, shifts the focus on the vocabulary’s natural domain. Through Entropy-Lens, we use information theory to study the evolution of entropy of the intermediate layers’ decoded logits.
@@ 3.1 Information Theory
Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’. These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible event $A$ and $B$. The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined.
Next, we study the entropy of vocabulary predictions—a quantity that is maximal when the prediction assigns equal probability to all tokens, minimal when it assigns zero probability to all but one token, and takes intermediate values when probability is distributed across multiple tokens, consistent with the previously mentioned properties.
@@ 3.2 The Transformer
The transformer (Vaswani et al., 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021). Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 2. While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich, 2019), attention, and multi layer perceptrons (MLPs), as shown in Fig. 1. Residual connections, instead, sum the output of the layer $i-1$ to the output of the layer $i$.
@@
LLMs are trained to simply predict the next token in a sentence. That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$. Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, sometimes set to $E^{\top}$, tying the embedding weights and the decoder. Finally, the logits—the output of $D$—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary. We summarize this operation with the function $W:=\operatorname{softmax}\circ\,D$.
@@
3.2.1 Instruct Models
Training a Large Language Model (LLM) requires vast amounts of data and is generally divided into multiple phases.
@@ 3.2.1 Instruct Models
For our experiments, we used off-the-shelf models without RLHF to analyze information processing in a less biased transformer version. We also focused on Instruct models instead of Chat models for two reasons: 1. the Instruct strategy aligns better with our experimental setup 2. Instruct models are more flexible and often preferred for practical applications.
@@ 4 Method
The aim of our framework is to find and characterize the information-theoretic signature of transformer computations. Entropy-Lens’s pipeline comprises three steps and is described in Figure 2.
@@
Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ defined in Definition 1. In our experiments, we set both of them to be concatenation, so that $h_{[N:T]}$ is the matrix with $h_{N+t}$ as columns, that is $(h_{[N:T]})_{t}^{i}=h_{N+t}^{i}$ for $i\in\{1,\dots,L\}$ and $t\in\{0,\dots,T-1\}$. The aggregation of entropy profiles is the step 2 of our framework.
The last step of our framework is classification, where we feed the aggregated entropy profile to a classifier $\mathcal{C}$ to determine whether it contains sufficient information to identify a particular ‘entity’.
In our experiments, we examine whether aggregated entropy profiles identify model family (Section 5.1.1), task type (Section 5.1.2), and correct and wrong answers to multiple choice questions (Section 5.1.3) in LLMs. Additionally, we apply the same pipeline to ViTs showing the flexibility of the proposed methodology (Section 5.2). In our experiments, we take $\mathcal{C}$ to be a k-NN classifier. Classification of the aggregated entropy profiles is the step 3 our framework.
@@ 5 Experiments
The goal of our experiments is to demonstrate the effectiveness of entropy profiles in distinguishing computational signatures across various models and scenarios. In the first part, we explore how entropy profiles can differentiate between LLM families and characterize distinct tasks. In the second part, we apply our framework to the Vision Transformer architecture to assess whether insights from the LLM experiments can be similarly recovered.
5.1 Entropy-Lens for Large Language Models
The experiments on LLMs focus on three key aspects. First, we demonstrate that entropy profiles are indicative of model families, with distinctions becoming more pronounced as model size increases. Second, we investigate whether the entropy profile of a model alone can be used to classify the task it is performing. Finally, we extend this analysis to assess whether entropy profiles can distinguish between correct and incorrect task execution.
5.1.1 Entropy profiles identify model families
We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 9 models from 3 different families (GPT, Gemma and LLama) with parameter counts ranging from 100M to 9B.
We compute the mean of the entropy profiles obtained from 64 generated tokens, obtained with the prompt ‘The concept of entropy, a brief essay:’, as shown in Figure 3. We observe that the profiles relate uniquely to the model family, rather than a particular model, independently of its size.
@@ 5.1.1 Entropy profiles identify model families
The Gemma model class, on the other hand, starts with low entropy in the very first layers, then rises to higher entropy in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
The Llama model class follows a similar pattern, but with a steeper rise, resulting in a higher entropy value maintained over a larger range of intermediate layers.
@@
Figure 3: Average entropy profiles over 64 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
5.1.2 Entropy profiles identify task types
@@
We do this with the TinyStories dataset (Eldan and Li, 2023). For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix A Table 3, and a story from TinyStories. These templates are:
@@
We generate 800 prompts per task type, $1/3$ of them with the base template, $1/3$ with the reversed template, and $1/3$ with the scrambled templates, for a total of 2400 prompts. We then apply our pipeline from Section 4 to classify the aggregated entropy profiles of these prompts against their task type using a k-NN classifier. The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest). Table 1 shows the results obtained for 6 models with parameter counts ranging from 1B to 9B. Figure 4 shows the average entropy profiles per task type.
@@
Figure 4: Average entropy profiles with shaded standard deviation for different task types: generative (continuation_prompt 1 and 2), syntactic (counting_prompt 1 and 2), and semantic (semantic_prompt 1 and 2). These tasks are induced with the prompts described in Appendix A. Left: Llama-3.2-Insruct. Right: Gemma-2-Instruct
Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. The standard deviation is calculated on the 10-fold cross-validation splits.
@@
Llama-3.2-Instruct 1 B 94.94 ± 0.79
Llama-3.2-Instruct 3 B 94.77 ± 0.93
Llama-3-Instruct 8 B 96.10 ± 0.67
@@
5.1.3 Entropy profiles identify correct task execution
We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset Hendrycks et al. (2021).
MMLU consists of multiple-choice questions across 57 subjects, ranging from history and physics to law, mathematics, and medicine. The difficulty levels span from elementary to professional, making it a benchmark for evaluating language models on specialized domains. Each dataset entry contains: a question string, four answer choices and a label indicating the correct answer.
We evaluated two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 4 in Appendix A):

### v2

@@ Abstract
Transformer models map input token sequences to output token distributions, layer by layer. While most interpretability work focuses on internal latent representations, we study the evolution of these token-level distributions directly in vocabulary space. However, such distributions are high-dimensional and defined on an unordered support, making common descriptors like moments or cumulants ill-suited. We address this by computing the Shannon entropy of each intermediate predicted distribution, yielding one interpretable scalar per layer. The resulting sequence—the entropy profile—serves as a compact, information-theoretic signature of the model’s computation.
We introduce Entropy-Lens, a model-agnostic framework that extracts entropy profiles from frozen, off-the-shelf transformers. We show that these profiles (i) reveal family-specific computation patterns invariant under depth rescaling, (ii) are predictive of prompt type and task format, and (iii) correlate with output correctness. We further show that Rényi entropies yield similar results within a broad range of $\alpha$ values, justifying the use of Shannon entropy as a stable and principled summary. Our results hold across different transformers, without requiring gradients, fine-tuning, or access to model internals.
00footnotetext: Preprint - This manuscript is currently under review.11footnotetext: Code available at: github.com/christopher-irw/Entropy-Lens
@@ 1 Introduction
Transformer-based architectures (Vaswani et al. 2023) operate by iteratively mapping token inputs to distributions over next-token outputs, layer by layer (Shan et al. 2024). This view emphasizes their inherently probabilistic nature: each layer produces a distribution over the vocabulary, encoding the model’s beliefs at that stage.
Despite their success across domains—from language to biology to vision (Devlin 2018; Ji et al. 2021; Wu et al. 2020)—the internal evolution of these distributions remains poorly understood, resulting in unpredictable behaviour (Wei et al. 2022) and reliability concerns (Schroeder and Wood-Doughty 2025; Huang et al. 2025). Most interpretability research either focuses on toy models (Elhage et al. 2021, 2022) or requires training a set of probes (nostalgebraist 2020; Belrose et al. 2023) or full models on ad-hoc tasks (Nanda et al. 2023), limiting scalability and generality. Moreover, existing methods primarily study latent representations—hidden states, attention maps, or MLP activations (Nanda and Bloom 2022; Bereska and Gavves 2024; Chefer, Gur, and Wolf 2021)—rather than the output distributions themselves.
Figure 1: TSNE of the aggregated entropy profiles of different model families using unprompted models. The size of each point is scaled based on the number of parameters of the model.
We propose to shift perspective: instead of analyzing the internal geometry of latent activations, we focus on the evolution of the model’s token-level predictions across layers. These predictions, obtained by decoding intermediate representations via the model’s output head, form probability distributions over the vocabulary. Studying their evolution offers a direct, vocabulary-grounded view of the computation.
However, these distributions pose two challenges: they are high-dimensional (one value per token) and defined over an unordered support (the vocabulary). This makes classical descriptors like variance or higher-order cumulants unstable or ill-defined. We address both issues by computing the Shannon entropy of each distribution—a scalar, interpretable quantity invariant to token permutation, and reflective of the model’s uncertainty at each layer. We introduce Entropy-Lens, a simple, scalable, and model-agnostic framework for studying transformer computations through the evolution of entropy across layers.
In our experiments, we consider several LLMs, including Llama (Touvron et al. 2023), Gemma (Team et al. 2024), and GPT, (Radford 2018) up to 9B parameters. For each generated token, we compute the entropy of its intermediate predictions, as described in Section 4, yelding one interpretable value per layer: this constitutes the entropy profile. We then aggregate these profiles across tokens and study whether they capture structural information about the internal LLM’s computation (Figure 1). In Section 5, we show that: (i) entropy profiles reveal family-specific signatures of transformer architectures, invariant under depth rescaling; (ii) they are predictive of prompt semantics and format; (iii) they correlate with output correctness; (iv) and results are stable across Rényi entropy variants, justifying our use of Shannon entropy as a principled default.
Entropy-Lens requires no gradients, no fine-tuning, and no access to internal weights. It applies to frozen, off-the-shelf transformers of arbitrary size. While our experiments focus on language models, the methodology is general and opens new directions for vocabulary-grounded interpretability.
@@ 2 Related Work
Mechanistic interpretability (Bereska and Gavves 2024) aims to provide a precise description and prediction of transformer-based computations. Common tools in the field are lenses, which are a broad class of probes deployed in intermediate steps of the residual stream. For example, logit-lens (nostalgebraist 2020) uses the model’s decoder function to decode the intermediate activations in the vocabulary space. tuned-lens (Belrose et al. 2023) refines this technique by training a different affine probe at each layer, instead of only using the pretrained model’s decoder function. Building on the Transformer-Lens library (Nanda and Bloom 2022), we propose Entropy-Lens, which employs logit-lens to study and characterize LLMs’ computations via their decoded version with information theory.
@@
Information theory has been studied both in connection to the training phase of LLMs and their interpretability. For example, a collapse in attention entropy has been linked to training instabilities (Zhai et al. 2023) and matrix entropy was employed to evaluate “compression” in LLMs (Wei et al. 2024). Additionally, mutual information was used to study the effectiveness of the chain-of-thought mechanism (Ton, Taufiq, and Liu 2024). Our work, instead, shifts the focus on the vocabulary’s natural domain. Through Entropy-Lens, we use information theory to study the evolution of entropy of the intermediate layers’ decoded logits. A related study by Dombrowski and Corlouer (2024) uses information-theoretic measures to distinguish between truthful and deceptive LLM generations. Their analysis focuses on deception under explicit instruction, whereas our work aims at uncovering broader entropy-based signatures across model families, prompt types, and output correctness without requiring behavioral framing or fine-tuned prompting.
@@ 3.1 Information theory
Shannon proved that this function is the only one—up to a scalar multiplication—that satisfies intuitive properties for measuring ‘disorder’ (Shannon 1948). These include being maximal for a uniform distribution, minimal for the limit of a Kronecker delta function, and ensuring that $H(A,B)\leq H(A)+H(B)$ for every possible event $A$ and $B$. The same function already existed in continuous form in physics, where it linked the probabilistic formalism of statistical mechanics with the more phenomenological framework of thermodynamics, where the term ‘entropy’ was originally coined (Gibbs 1902). In addition to Shannon entropy, we also consider its generalization known as Rényi entropy. Given a discrete random variable $X$ with probability mass function $p$, the Rényi entropy of order $\alpha>0$, $\alpha\neq 1$, is defined as:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\sum_{i}p(x_{i})^{\alpha}$ (2)
This formulation reduces to Shannon entropy in the limit $\alpha\to 1$, and introduces a tunable parameter $\alpha$ that modulates the sensitivity of the entropy to the distribution’s tail. Lower values of $\alpha$ give more weight to rare events, while higher values emphasize the most probable outcomes.
Next, we study the entropy of vocabulary predictions—a quantity that is maximal when the prediction assigns equal probability to all tokens, minimal when it assigns zero probability to all but one token, and takes intermediate values when probability is distributed across multiple tokens, consistent with the previously mentioned properties. In our experiments, we explored a range of $\alpha$ values to understand how this parameter affects the informativeness of the resulting entropy profiles. We identified an informative range of $\alpha$ values—which includes the Shannon case $\alpha=1$—where entropy profiles retain high discriminative power for classification tasks. Outside of this range, we observe that profiles tend to collapse: for very small $\alpha$, entropies become nearly maximal and lose contrast across layers and examples; for large $\alpha$, they become very small and overly sensitive to noise and local fluctuations. These findings support the use of Shannon entropy as a balanced and parameter-free choice, robust across a wide range of practical conditions.
@@ 3.2 The Transformer
The transformer (Vaswani et al. 2023) is a deep learning architecture widely applied in language modelling with LLMs (Brown et al. 2020) and computer vision (Dosovitskiy et al. 2021). Transformer computations happen through transformer blocks and residual connections, as exemplified in Figure 3. While various design choices are possible, blocks are usually a composition of layer normalization (Zhang and Sennrich 2019), attention, and multi layer perceptrons (MLPs), as shown in Figure 2. Residual connections, instead, sum the output of the layer i1i-1 to the output of the layer $i$.
@@
LLMs are trained to predict the next token in a sentence. That is, given a sentence prompt $S$ with tokens $t_{1},\dots,t_{N}$, the transformer encodes each token with a linear encoder $E$. Throughout the residual stream, the representation $x_{N}$ of the token $t_{N}$ evolves into the representation of the token $t_{N+1}$, which is then decoded back into token space via a linear decoder $D$, often set to $E^{\top}$, tying the embedding weights and the decoder. Finally, the logits—the output of DD—are normalized with $\operatorname{softmax}$ to represent a probability distribution over the vocabulary. We summarize this operation with the function $W:=\operatorname{softmax}\circ\,D$.
@@
Instruct models
Training an LLM requires vast amounts of data and is generally divided into multiple phases.
@@
For our experiments, we used off-the-shelf models without RLHF to analyze information processing in a less biased transformer version. We also focused on Instruct models (abbreviated with it in tables and figures) instead of Chat models for two reasons: 1. the Instruct strategy aligns better with our experimental setup 2. Instruct models are more flexible and often preferred for practical applications.
@@ 4 Method
Entropy-Lens’s pipeline comprises three steps and is described in Figure 3. After introducing the notation and motivating the choice of entropy as a measure, in the following we describe the details of the framework which as been used for the experiments in the following sections.
@@
Why entropy?
Our goal is to characterize how a transformer model evolves its predictions across layers, remaining in the token space for interpretability. Each intermediate representation, once decoded and passed through softmax, yields a probability distribution over the vocabulary. However, this distribution lives on a high-dimensional, unordered categorical support.
Classical descriptors such as variance or higher-order cumulants rely on an implicit ordering of the support and thus become meaningless when applied to token distributions—shuffling token indices alters their value arbitrarily. Entropy, on the other hand, is invariant under permutations and captures a well-defined notion of uncertainty or informativeness regardless of vocabulary indexing. It is therefore a natural and stable choice to summarize the evolution of token-level beliefs across transformer layers.
Definitions
@@
Note that $\bigotimes$ in Definition 2 need not be the same as $\bigoplus$ in Definition 1. In our experiments, we set both of them to be concatenation, so that $h_{[N:T]}$ is the matrix with $h_{N+t}$ as columns, that is $(h_{[N:T]})_{t}^{i}=h_{N+t}^{i}$ for $i\in\{1,\dots,L\}$ and $t\in\{0,\dots,T-1\}$. The aggregation of entropy profiles is the step 2 of our framework.
The last step of our framework is classification, where we feed the aggregated entropy profile to a classifier $\mathcal{C}$ to determine whether it contains sufficient information to identify a particular ‘entity’. In our experiments, we take $\mathcal{C}$ to be a k-NN classifier.
We also examine whether aggregated entropy profiles identify model family (Section 5.1), task type and format (Sections 5.2 and 5.5), and correct and wrong answers to multiple choice questions (Section 5.3) in LLMs.
@@ 5 Experiments
Our experiments focus on several key aspects. First, we show that entropy profiles are indicative of model family, with distinctions becoming more pronounced as model size increases. We then investigate whether a model’s entropy profile alone can be used to classify the task it is performing. Next, we assess whether entropy profiles capture information about output formatting, independently of task content. We extend this analysis to evaluate whether entropy profiles can distinguish between correct and incorrect task execution. Additionally, we explore the effect of varying the alpha parameter in Rényi entropy, examining how it influences the expressiveness of the resulting profiles.
5.1 Entropy profiles identify model families
We assess whether aggregated entropy profiles can distinguish different model families by visualizing and analyzing those of 12 models from 4 different families (GPT, Gemma, LLama and Qwen) with parameter counts ranging from 100M to 9B.
We average 64 entropy profiles obtained by generating 32 tokens prompting the model with a completely blank prompt. We observe (see Figure 1) that the profiles relate uniquely to the model family, rather than a particular model, independently of its size. Moreover, we observe that each model size within a particular family is tied to a scaling factor if we normalize by number of layers (see Figure 4).
@@ 5.1 Entropy profiles identify model families
The Gemma model class, on the other hand, starts with high entropy in the very first layer, then sharply drops to lower entropy, rises again in the intermediate layers, and finally decreases to low entropy again just before the last layers, where the model is required to produce an output token.
The Llama model class starts with low entropy , then steeply rises and maintains a high entropy value over a large range of intermediate layers, finally decreasing to low entropy again.
The Qwen model family exhibits a similar trend, but in a more gradual manner, resulting in less well-defined regimes.
@@
Figure 4: Average of 64 entropy profiles over 32 generated tokens per model. The $x$ axis is normalized for an easier comparison when models have a different number of layers.
5.2 Entropy profiles identify task types
@@
We do this with the TinyStories dataset (Eldan and Li 2023). For evaluation robustness, we construct for each task type three prompt templates using a combination of task-specific task prompts, reported in Appendix C.1 Table 4, and a story from TinyStories. These templates are:
@@
We generate 800 prompts per task type, $1/3$ of them with the base template, $1/3$ with the reversed template, and $1/3$ with the scrambled templates, for a total of 2400 prompts. We then apply our pipeline from Section 4 to classify the aggregated entropy profiles of these prompts against their task type using a k-NN classifier. The model was evaluated in a 10-fold cross-validation using the ROC-AUC score (one-vs-rest), a standard choice for measuring classification performance. Table 1 shows the results obtained for 6 models with parameter counts ranging from 1B to 9B. Figure 5 shows the average entropy profiles per task type.
@@
Figure 5: Average entropy profiles with shaded standard deviation for different task types: generative, syntactic, and semantic. These tasks are induced with the prompts described in Appendix C.1. Left: Llama-3.2-it. Right: Gemma-2-it.
@@
Llama-3.2-it 1 B 94.94 ± 0.79
Llama-3.2-it 3 B 94.77 ± 0.93
Llama-3-it 8 B 96.10 ± 0.67
@@
Table 1: ROC-AUC (one-vs-rest) results of different models on the TinyStories task classification. Standard deviations are computed over 10-fold cross-validation splits.
5.3 Entropy profiles identify correct task execution
We test whether entropy profiles can identify correct and wrong answers generated by LLMs using the Massive Multitask Language Understanding (MMLU) dataset (Hendrycks et al. 2021). MMLU consists of multiple-choice questions across 57 subjects, ranging from history and physics to law, mathematics, and medicine. The difficulty levels span from elementary to professional, making it a benchmark for evaluating language models on specialized domains. Each dataset entry contains: a question string, four answer choices and a label indicating the correct answer.
We evaluate two models, a Llama-3.2 with 3B parameters Instruct and a Gemma-2 with 2B parameters, by presenting the multiple-choice questions in three different formats (as per Table 5 in Appendix C.2):